# CTRL-F: Pairing Convolution with Transformer for Image Classification via Multi-Level Feature Cross-Attention and Representation Learning Fusion

### v1

@@ 4.3 Main Results
We seek to demonstrate the effectiveness of CTRL-F variants, evaluating their performance on both a small and challenging dataset (i.e., Oxford-102 Flowers dataset), and on a diverse and large dataset (i.e., PlantVillage dataset). We compare the performance of our CTRL-F variants against some of the current state-of-the-art methods with comparable computational cost (i.e., parameters and FLOPs) belonging to the categories of pure ConvNets, Vision Transformers (ViTs), and Hybrid models. To ensure a fair comparison, all models are trained from scratch (i.e., without any pre-training) on both datasets using an input size of 224224. A state-of-the-art comparison in terms of top-1 accuracy, number of parameters, and FLOPs is given in Table 7.
@@
Comparison with CNNs. We compare our proposed variants against some state-of-the-art CNN-based models with comparable Params and FLOPs. Our CTRLF-S (+AKF) obtains 80.79% top-1 accuracy, surpassing lightweight CNNs including MobileNetV3-L and EfficientNet-B3 on Oxford-102 dataset with an absolute margin of 17.4% and 11.7% respectively, and our CTRLF-S (+CKF) obtains 99.91% on PlantVillage dataset, surpassing MobileNetV3-L and EfficientNet-B3 with a margin of 0.12% and 0.06%, respectively. Further our larger variant CTRLF-B (+AKF) surpasses ConvNeXt-T and EfficientNetV2-S with a margin of 30.3% and 19.7% on Oxford-102 dataset respectively, and our CTRLF-B (+CKF) surpasses ConvNeXt-T and EfficientNetV2-S with a margin of 0.32% and 0.07% on PlantVillage dataset, respectively. It can be deduced that our proposed variants significantly outperform the widely used CNNs in terms of accuracy with comparable cost. Also, this reveals the efficiency of our proposed variants when trained from scratch and without any extra data on a very small dataset (i.e., Oxford-102) and on a larger dataset (i.e., PlantVillage).
@@
Comparison with transformers. Due to their reliance on extensive datasets, pure transformer models often require large amounts of data to achieve state-of-the-art results. Consequently, when trained from scratch on small datasets, the performance of these transformers typically falls behind fully convolutional networks and hybrid models. Our variants significantly outperform recent ViTs. The CTRLF-S (+AKF) variant surpasses CrossViT-15, PVTv1-S and Swin-T with an absolute margin of 32.2%, 27.9% and 20.0% on the Oxford-102 dataset respectively, and our CTRLF-S (+CKF) surpasses CrossViT-15, PVTv1-S and Swin-T by margins of 0.54%, 0.32% and 0.20% on PlantVillage dataset while being smaller and more resource efficient. Additionally, our CTRLF-B (+AKF) variant surpasses CrossViT-15, PVT-v1-S and Swin-T by 33.6%, 29.4% and 21.4% on the Oxford-102 dataset respectively, and CTRLF-B (+CKF) surpasses CrossViT-15, PVT-v1-S and Swin-T by 0.54%, 0.32% and 0.20% on the PlantVillage dataset, respectively.
@@
Comparison with hybrid models. Recent advancements in hybrid approaches succeeded to leverage the complementary capabilities of Convolutional Neural Networks (ConvNets) and Vision Transformers (ViTs), surpassing their individual capabilities across various benchmark datasets such as ImageNet. We trained recent SOTA hybrid models from scratch on Oxford-102 and PlantVillage dataset, comparing our proposed hybrid variants against them. Our proposed hybrid variants showed favorable performance in terms of top-1 accuracy on both datasets compared to other experimented hybrid models. For instance, our CTRLF-S (+AKF) significantly outperforms PVTv2-B1 by 22.4% on Oxford-102 dataset with a slight decrease in Params and FLOPs and our CTRLF-S (+CKF) outperforms PVTv2-B1 by 0.16% on PlantVillage dataset. Meanwhile, our CTRLF-B (+AKF) outperforms PVTv2-B2, Coat-0, Coat-1 and Moat-0 by 22.0%, 11.2%, 10.5% and 5.8% on Oxford-102 dataset respectively, with a slight decrease in Params and FLOPS, and on PlantVillage dataset our CTRLF-B (+CKF) surpasses PVTv2-B2, Coat-0, Coat-1 and Moat-0 by 0.18%, 0.06%, 0.06% and 0.07%, respectively.
Model | Params | FLOPs | Oxford-102 | PlantVillage | |
top-1 acc | top-acc | ||||
ConvNets | MobileNetV3-L | 4.3M | 0.22G | 63.42% | 99.79% |
EfficientNet-B3 | 10.9M | 1.0G | 69.10% | 99.85% | |
EfficientNetV2-S | 20.3M | 2.9G | 62.52% | 99.84% | |
ConvNeXt-T | 27.9M | 4.5G | 51.96% | 99.59% | |
ViTs | ViT-B | 85.9M | 17.6G | 46.50% | 98.74% |
CrossViT-15† | 27.7M | 6.1G | 48.61% | 99.37% | |
PVTv1-S | 24.0M | 3.7G | 52.85% | 99.59% | |
Swin-T | 27.6M | 4.5G | 60.82% | 99.71% | |
Hybrid | PVTv2-B1 | 13.5M | 2.1G | 58.40% | 99.75% |
PVTv2-B2 | 24.9M | 3.9G | 60.22% | 99.73% | |
Coat-0 | 25.0M | 4.2G | 71.00% | 99.85% | |
Coat-1 | 41.0M | 8.4G | 71.76% | 99.85% | |
Moat-0 | 27.0M | 5.7G | 76.46% | 99.84% | |
CTRLF-S (+AKF) | 9.97M | 1.43G | 80.79% | 99.85% | |
CTRLF-B (+AKF) | 21.36M | 3.29G | 82.24% | 99.89% | |
CTRLF-S (+CKF) | 9.99M | 1.43G | 79.23% | 99.91% | |
CTRLF-B (+CKF) | 21.39M | 3.29G | 81.46% | 99.91% |
@@ 5 Conclusion
In this paper, we introduce CTRL-F, a novel lightweight hybrid architecture that seamlessly integrates the strengths of Convolutional Neural Networks (ConvNets) and Vision Transformers (ViTs) for image classification tasks. To effectively capture global context with reduced complexity, we introduce a Multi-level Feature Cross-Attention Transformer (MFCA) module with a dual-branch structure. This module efficiently operates on low-resolution, multi-level feature maps obtained from two different convolutional stages and process these feature maps independently using a series of transformer encoders and facilitates efficient knowledge exchange between both branches via cross-attention. We also develop two effective strategies for representation learning fusion named: adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF), providing complementary benefits by ensuring that the model incorporates both local representations learned from convolution and global representations learned from attention. Extensive experiments demonstrate that our proposed CTRL-F enjoys ConvNets benefits including fast convergence and robust generalization with limited data, alongside the high capacity and global processing capabilities of transformer, achieving state-of-the-the-art results on different datasets with compact computational cost.

### v2

@@ 4.3 Main Results
We seek to demonstrate the effectiveness of CTRL-F variants, evaluating their performance on both a small and challenging datasets (i.e., Oxford-102 Flowers and CUB-200 datasets), and on a diverse and large dataset (i.e., PlantVillage dataset). We compare the performance of our CTRL-F variants against some of the current state-of-the-art methods with comparable computational cost (i.e., parameters and FLOPs) belonging to the categories of pure ConvNets, Vision Transformers (ViTs), and Hybrid models. To ensure a fair comparison, all models are trained from scratch (i.e., without any pre-training) on all datasets using an input size of 224224. A state-of-the-art comparison in terms of top-1 accuracy, number of parameters, and FLOPs is given in Table 7.
@@
Comparison with CNNs. We compare our proposed variants against some state-of-the-art CNN-based models with comparable Params and FLOPs. Our CTRLF-S (+AKF) surpasses lightweight CNNs including MobileNetV3-L and EfficientNet-B3 with an absolute margin of 17.0% and 12.6% respectively, on the Oxford-102 dataset. Our larger variant, CTRLF-B (+AKF) outperforms ConvNeXt-T and EfficientNetV2-S by 30.6% and 20.6% respectively, on Oxford-102 dataset. Similarly, CTRLF-S (+CKF) surpasses MobileNetV3-L and EfficientNet-B3 by 17.9% and 9.3%, while CTRLF-B (+CKF) surpasses ConvNeXt-T and EfficientNetV2-S by 29.9% and 27.6% on CUB-200 dataset. On the PlantVillage dataset, CTRLF-B (+CKF) achieves 0.32% and 0.07% higher accuracy than ConvNeXt-T and EfficientNetV2-S, respectively. It can be deduced that our proposed variants significantly outperform the widely used CNNs in terms of accuracy with comparable cost. Also, this reveals the efficiency of our proposed variants when trained from scratch and without any extra data on both small datasets (Oxford-102 and CUB-200) and on a larger dataset (PlantVillage).
@@
Comparison with transformers. Due to their reliance on extensive datasets, pure transformer models often require large amounts of data to achieve state-of-the-art results. Consequently, when trained from scratch on small datasets, the performance of these transformers typically falls behind pure CNNs and hybrid models. Our variants significantly outperform recent ViTs. The CTRLF-S (+AKF) variant surpasses CrossViT-15, PVTv1-S and Swin-T with an absolute margin of 33.8%, 29.6% and 21.6% on the Oxford-102 dataset respectively, and our CTRLF-S (+CKF) surpasses CrossViT-15, PVTv1-S and Swin-T by margins of 0.54%, 0.32% and 0.20% on PlantVillage dataset and by 36.6%, 39.4% and 28.8% on CUB-200 dataset while being smaller and more resource efficient. Additionally, our CTRLF-B (+AKF) variant surpasses CrossViT-15, PVTv1-S and Swin-T by 35.4%, 31.2% and 23.2% on the Oxford-102 respectively, and CTRLF-B (+CKF) outperforms CrossViT-15, PVTv1-S and Swin-T by 0.54%, 0.32% and 0.20% on the PlantVillage and by 40.9%, 43.7% and 33.1% on CUB-200.
@@
Comparison with hybrid models. Recent advancements in hybrid approaches have successfully leveraged the complementary capabilities of Convolutional Neural Networks (ConvNets) and Vision Transformers (ViTs), surpassing their individual capabilities across various benchmark datasets such as ImageNet. We trained recent SOTA hybrid models from scratch on Oxford-102, CUB-200 and PlantVillage datasets, comparing our proposed hybrid variants against them. Our proposed hybrid variants showed favorable performance in terms of top-1 accuracy compared to other evaluated hybrid models. For instance, our CTRLF-S (+AKF) significantly outperforms PVTv2-B1 by 24.0% on Oxford-102 dataset with a slight decrease in Params and FLOPs and our CTRLF-S (+CKF) outperforms PVTv2-B1 by 25.7% on CUB-200 and by 0.16% on PlantVillage dataset. Meanwhile, our CTRLF-B (+AKF) outperforms PVTv2-B2, FasterViT-0 ,Coat-0, Coat-1 and Moat-0 by 23.8%, 19.7%, 12.9%, 12.2% and 7.4% on Oxford-102, with a slight decrease in Params and FLOPS, and our CTRLF-B (+CKF) surpasses PVTv2-B2, FasterViT-0, Coat-0, Coat-1 and Moat-0 by 28.2%, 22.0%, 7.5%, 4.0% and 22.3% on CUB-200, and by 0.18%, 0.15%, 0.06%, 0.06% and 0.07% on PlantVillage.
| Model | Params | FLOPs | Oxford-102 | CUB-200 | PlantVillage | |
| top-1 acc | top-1 acc | top-1 acc | ||||
| ConvNets | MobileNetV3-L [40] | 4.3M | 0.22G | 65.46% | 39.42% | 99.79% |
| EfficientNet-B3 [43] | 10.9M | 1.0G | 69.85% | 48.05% | 99.85% | |
| EfficientNetV2-S [20] | 20.3M | 2.9G | 63.36% | 34.09% | 99.84% | |
| ConvNeXt-T [21] | 27.9M | 4.5G | 53.40% | 31.81% | 99.59% | |
| ViTs | ViT-B [9] | 85.9M | 17.6G | 46.50% | 12.70% | 98.74% |
| CrossViT-15†[17] | 27.7M | 6.1G | 48.61% | 20.75% | 99.37% | |
| PVTv1-S [16] | 24.0M | 3.7G | 52.85% | 17.93% | 99.59% | |
| Swin-T [18] | 27.6M | 4.5G | 60.82% | 28.55% | 99.71% | |
| Hybrid | PVTv2-B1 [55] | 13.5M | 2.1G | 58.40% | 31.62% | 99.75% |
| PVTv2-B2 [55] | 24.9M | 3.9G | 60.22% | 33.47% | 99.73% | |
| FasterViT-0 [56] | 30.9M | 3.3G | 64.31% | 39.70% | 99.76% | |
| Coat-0 [28] | 25.0M | 4.2G | 71.10% | 54.14% | 99.85% | |
| Coat-1 [28] | 41.0M | 8.4G | 71.81% | 57.66% | 99.85% | |
| Moat-0 [29] | 27.0M | 5.7G | 76.57% | 39.42% | 99.84% | |
| CTRLF-S (+AKF) | 9.97M | 1.43G | 82.42% | 56.27% | 99.85% | |
| CTRLF-B (+AKF) | 21.36M | 3.29G | 84.01% | 60.77% | 99.89% | |
| CTRLF-S (+CKF) | 9.99M | 1.43G | 81.88% | 57.35% | 99.91% | |
| CTRLF-B (+CKF) | 21.39M | 3.29G | 83.00% | 61.68% | 99.91% |
@@ 5 Conclusion
In this paper, we introduce CTRL-F, a novel lightweight hybrid architecture that seamlessly integrates the strengths of Convolutional Neural Networks (ConvNets) and Vision Transformers (ViTs) for image classification tasks. To effectively capture global context with reduced complexity, we introduce a Multi-level Feature Cross-Attention Transformer (MFCA) module with a dual-branch structure. This module efficiently operates on low-resolution, multi-level feature maps obtained from two different convolutional stages and process these feature maps independently using a series of transformer encoders and facilitates efficient knowledge exchange between both branches via cross-attention. We also develop two effective strategies for representation learning fusion named: adaptive knowledge fusion (AKF) and collaborative knowledge fusion (CKF), providing complementary benefits by ensuring that the model incorporates both local representations learned from convolution and global representations learned from attention. Extensive experiments demonstrate that our proposed CTRL-F enjoys ConvNets benefits including fast convergence and robust generalization with limited data, alongside the high capacity and global processing capabilities of transformer, achieving state-of-the-art results on different datasets with compact computational cost.